OpenStudy (openstudyoperator):

Which expresses 45:50 as a decimal? Question 10 options: A) 45.50 B) 0.9 C) 9.10 D) 50.45

OpenStudy (openstudyoperator):

@akiaki28

OpenStudy (skullpatrol):

any ideas?

OpenStudy (skullpatrol):

$$\Huge a:b =a \div b = \dfrac{a}{b}$$

OpenStudy (openstudyoperator):

Uhhhhh

OpenStudy (skullpatrol):

$$\Huge 45:50=45 \div50 =\dfrac{45}{50}$$

OpenStudy (skullpatrol):

$$\Huge \dfrac{45}{50}=\dfrac{9 \cdot 5}{10 \cdot 5}=\dfrac{9 \cdot \cancel 5}{10 \cdot \cancel 5}=\dfrac{9}{10} $$

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):